What is the magnitude of the gravitational field (in pN/kg) from a uniformly dense square plane at a position a distance d=2m away from one side of the plane, along the perpendicular bisector of the plane? The net mass of the plane is M=10.5 kg and the length of each side of the plane is L=0.85 m.
Answer #1

Distance, r = (0.85/2) + 2 = 2.425 m

Gravitational field, g = GM/r^2

g = (6.67 x 10^-11 x 10.5)/(2.425^2) = 1.19 x 10^-10 = 119.09 pN/kg

Comment in case any doubt please rate my answer...